A ray of light is incident on the surface of a crown glass plate at an angle of 30 with the normal. Determine the angle between the refracted ray and reflected ray.

1
Expert's answer
2022-05-18T15:27:57-0400

The Snell's law says

"n_i\sin\theta_i=n_t\sin\theta_t"

"\theta_t=\sin^{-1}\frac{n_i\sin\theta_i}{n_t}"

"\theta_t=\sin^{-1}\frac{1.0*\sin30^\circ}{1.52}=19^\circ"

"\theta_r=\theta_i=30^\circ"

The angle between the refracted ray and reflected ray is given by

"\theta=180^\circ-(\theta_r+\theta_t)\\
=180^\circ-(30^\circ+19^\circ)=131^\circ"